When a test charge of +4.00 nC is placed at a certain point, the force that acts on it has a magnitude of 0.0100 N and is directed northeast.

3)

What is the electric field at the point in question? (Express your answer to three significant figures.)

When a test charge of +4.00 nC is placed at a certain point, the force that acts on it has a magnitude of 0.0100 N and is directed northeast. 1) If the test charge were -2.00 nC instead, what would be the magnitude of the force that would act on it? (Express your answer to three significant figures.) 005 N Submit You currently have 6 submissions for this question. Only 10 submission are allowed You can make 4 more submissions for this question 2) What would be the direction of the force that would act on it? northeast northwest southwest southeast Submit You currently have 1 submissions for this question. Only 10 submission are allowed You can make 9 more submissions for this question 3) What is the electric field at the point in question? (Express your answer to three significant figures.) 05 x10 N/C Submit You currently have 5 submissions for this question. Only 10 submission are allowed You can make 5 more submissions for this question
Answer #1

Force is directly proportional to the magnitude of the charge,

F = qE where E is electric field and q is charge.

1) As the magnitude of the charge is halved (from 4 nC to 2 nC), the force will be halved,

Force = 0.01/2 = 0.005 N

2) Direction will be opposite to the positive charge's direction, that is southwest

3) Electric field = F/q = 0.01/4*10^-9 = 2500000 N = 0.25 * 10^7 N/C

Answer is 0.25 for 3rd question.